Some years ago, Penny purchased the car of her dreams for $25,000 by paying 20% down at purchase time and taking a $20,000, 5 years, 6% per year, compounded monthly loan with 60 monthly payments of $3866.66 each. She is examining her loan situation and would like to have some specific information. Help her obtain the following:

(a) Verification of the current monthly payment amount.

(b) Total amount she will pay over 5 years.

(c) Total interest she will pay over the 5 years and the percentage this represent of the original loan amount of $20,000.

(d) After she missed payment #36 at the very end of the third year, according to the loan agreement, the interest rate increased from 6% to 10% per year, compounded monthly. Based on the remaining principal immediately after the late payment, determine the new monthly payment. Verify that this increased amount is necessary to pay off the loan at the increased rate.

(e) Penny is now in her fourth year, has paid the increased payment for 12 payments, and wants to get rid of this loan completely. She wishes to know the remaining principal amount when payment #48 is due. There is no penalty for early repayment of principal.

Please put information in excel to get answers
Answer #1
(a) Verification of current monthly payment
Rate Monthly interest rate =(6/12)% 0.50%
Nper Number of months of payment=5*12 60
Pv Loan amount $20,000
PMT Monthly payment required $386.66 (Using PMT function of excel with Rate=0.5%, Nper=60, Pv=-20000)
(b) Total amount she will pay over 5 years $23,319.60 (388.66*60)
23319.60
.(c) Total interest payment =23319.60-20000 $3,319.60
Percentage of Interest of original loan 16.60% (3319.60/20000)
(d) Loan Balance after payment of 36th payment =Present Value of future payments
Number of Future Payments =60-36 24
Loan Balance after payment of 36th payment $8,724.07 (Using PV function of excel with Rate=0.5%, Nper=24, Pmt=-386.66)
Pv Loan Balance immediately after late payment $8,724.07
Rate Monthly interest rate =(10/12)% 0.83333%
Nper Number of Future Payments 24
PMT New Monthly payment required $402.57 (Using PMT function of excel with Rate=0.83333%, Nper=24, Pv=-8724.07)
.(e) REMAINING PRINCIPAL AMOUNT
Nper Number of future payment after 48th payment 12 (60-48)
Remaining Principal after payment of 48th payment $4,579.05 (Using PV function of excel with Rate=0.8333%, Nper=12, Pmt=-402.57)
Remaining Principal JUST BEFORE payment of 48th payment $4,981.62 (4579.05+402.57)
remaining principal amount when payment #48 is due $4,981.62

O G7 - X Fax =PMT(G4, G5,-G6) G H I J K L M N 2 3 Verification of current monthly payment 4 Monthly interest rate =(6/12)% 0.Luuuuu x fc =PV(0.5%,G16,-G7) G17 A B C D E G H I J K L M N 0.50% (a) Rate Nper Pv PMT Verification of current monthly paymenстролита G21 x fo =PMT(G19,620,-G18) C D E H I J K L M N G $20,000 $386.66 Loan amount Monthly payment required PMT (Using PMG25 fc =PV(G19,624,-G21) Дc D E J K L. M G 16.60% H I (3319.60/20000) Percentage of Interest of original loan (d) Loan Balanc